### v1

# LLM Roleplay: Simulating Human-Chatbot Interaction

Abstract
The development of chatbots requires collecting a large number of human-chatbot dialogues to reflect the breadth of users’ sociodemographic backgrounds and conversational goals. However, the resource requirements to conduct the respective user studies can be prohibitively high and often only allow for a narrow analysis of specific dialogue goals and participant demographics. In this paper, we propose LLM-Roleplay: a goal-oriented, persona-based method to automatically generate diverse multi-turn dialogues simulating human-chatbot interaction. LLM-Roleplay can be applied to generate dialogues with any type of chatbot and uses large language models (LLMs) to play the role of textually described personas. To validate our method we collect natural human-chatbot dialogues from different sociodemographic groups and conduct a human evaluation to compare real human-chatbot dialogues with our generated dialogues. We compare the abilities of state-of-the-art LLMs in embodying personas and holding a conversation and find that our method can simulate human-chatbot dialogues with a high indistinguishability rate.111https://github.com/UKPLab/llm-roleplay
T1zi4lmr
LLM Roleplay: Simulating Human-Chatbot Interaction
1 Introduction
Collecting human-chatbot dialogues requires recruiting and managing a large number of human annotators, which can pose prohibitive obstacles to researchers who aim to develop conversational AI agents (i.e., chatbots). To circumvent the latter and the limitations of publicly available data, numerous methods employing chatbots to generate dialogues have been introduced lately Xu et al. (2023b); Kim et al. (2023); Zhu et al. (2023); Ding et al. (2023); Zhao et al. (2024b). These methods have the potential to generate synthetic data more quickly and cost-effectively, offering a level of quality and variety that can closely match what human annotators can achieve Zhang et al. (2024).
Recently, several studies have leveraged such synthetic data generated through distillation and self-improvement techniques to enhance the capabilities of pre-trained large language models (LLMs) Zhang et al. (2024). By employing instruction tuning and fine-tuning techniques, these efforts aim to ensure that LLMs deliver useful and safe responses that adhere to provided instructions Xu et al. (2023a); Taori et al. (2023); Mukherjee et al. (2023); Li et al. (2023b); Peng et al. (2023); Almazrouei et al. (2023).
However, today’s dialogue generation methods have two critical limitations. First, the set of possible user responses for a given chatbot utterance is large while the existing datasets typically only cover a single or few annotated user responses. The previously generated datasets are constrained by the number of dialogues and turns, making it difficult to extend them to novel domains or conversational goals. Second, the existing methods rarely account for subjective response behavior and implicitly assume an average user, which can create a gap in representation Rottger et al. (2022) and poses the danger of evaluating and ultimately improving chatbots for a specific sociodemographic group while ignoring others.
The quality of existing datasets is primarily evaluated with the performance in a supervised fine-tuning (SFT) setup, which is influenced by various properties of the generated dataset. For example, Chen et al. (2023) argue that the initial prompt plays a crucial role in the quality of the generated dialogue. On the other hand, Zhao et al. (2024a) show that the length of the response is the key factor and with significantly fewer samples outperforms methods that solely focus on quality. Meanwhile, Shen (2024) emphasizes the importance of mimicking human-style interactions in the dialogues.
To mitigate the shortcomings of existing dialogue generation methods and to account for the desiderate identified in prior work for SFT, we introduce LLM Roleplay: a goal-oriented, persona-based, plug-and-play method to generate diverse, multi-turn, and long dialogues simulating human-chatbot interactions using LLMs. Our method is the first to instruct an LLM (inquirer) to mimic a given persona and prompt a chatbot (responder) to achieve a given conversational goal and thereby elicit realistic human-AI interactions with a specific LLM. Figure 1 shows an exemplary application of our method.
In this work, we address the following two research questions: (i) To what extent can we simulate real human-chatbot dialogues using LLM-chatbot dialogues? (ii) How do various LLMs compare as inquirers within this setup? To answer those research questions, we conduct two user studies. First, we collect real human-chatbot dialogues by asking participants to reach various conversational goals and link the collected dialogues with participants’ sociodemographic background. Next, we use LLM-Roleplay to simulate dialogues with the same set of personas and goals. Second, we conduct a human evaluation with another pool of participants to assess how well the generated dialogues mimic the collected ones. Specifically, we present two dialogues: one collected from the previous study and one simulated using LLM Roleplay, using the same persona and the conversation goal. We then ask the participants to try to identify the simulated one similar to a Turing test Turing (1950).
Overall, this paper contributes: (i) A novel method to simulate human-chatbot dialogues with an arbitrary set of personas and conversational goals, (ii) a human validation study that confirms our method’s potential to simulate human-chatbot interaction with high resemblance to real dialogues, (iii) a dataset of goal-oriented human-chatbot and model-chatbot dialogues using our hand-crafted multi-hop goals involving four state-of-the-art LLMs, (iv) an in-depth comparison of open-source and proprietary LLMs regarding their ability to embody persona and hold conversation, (v) an open-source implementation of our plug-and-play method that can be directly applied to simulate dialogues with any combination of models and chatbot systems using any set of conversational goals and personas.
2 Large Language Model Roleplay
In the following, we introduce the notation that we are following throughout this paper. Refer to Figure 1 for an annotated example.
Persona ()
is the composition of sociodemographic features. We conceptualize persona as a written description of an individual from a specific sociodemographic group. We follow Kumar et al. (2021) to define the sociodemographic features and the options of the latter.
Goal ()
is the textual representation of the conversational goal.
Subject ()
is a dialogue participant. We denote the inquirer as , the entity that asks questions, and the responder as , the entity that answers the given questions. In our setup, the inquirer () is an LLM and the responder () is a conversational agent or chatbot (not necessarily an LLM). The human inquirer will be denoted as .
Utterance ()
is the output of a Subject (), i.e. either the output of the inquirer () or the responder (). For example, will denote the inquirer’s -th utterance. We refer to two consecutive utterances of two different subjects as a turn: .
Dialogue ()
is a sequence of one or more turns. The dialogue of turns is denoted as: . We denote the maximum number of turns by , i.e. which we discuss in more detail in Section 3.1.
At a high level, LLM-Roleplay consists of three main steps: (i) Initial conditioning of the inquirer model with persona-specific text alongside the goal description. (ii) Subsequently, the inquirer’s output is provided to the responder model. (iii) Lastly, the output of the responder is returned to the inquirer by asking it to either output a follow-up question or a pre-defined token to terminate the dialogue.
We begin with assembling prompting templates for both of the subjects. We create a system prompt template (SYS_I) for the inquirer LLM. This template prompts the LLM to embody the given persona, provide a prompt to address the designated goal, and output the specified termination token if it considers the goal accomplished. As for the responder, we use a default system prompt (SYS_R) to promote it to be a helpful and honest assistant. Additionally, we develop a response forwarder template (INTER_I) for the inquirer. This template prompts the inquirer to assess the conclusiveness of the answer of the responder, determining whether to output a subsequent question or the termination token. We provide the system and forwarder prompt templates in Table 5 in Appendix B. The algorithm for the LLM Roleplay is shown in Algorithm 1.
To obtain a dialogue for a given persona and a goal we create a prompt by passing the persona () and the goal () to the inquirer system prompt template and generate an output based on inquirer LLM distribution:
A deterministic prompt extraction function, extract_prompt, that looks for a string in double quotes in the response, is applied to the latter to extract the prompt from the response:
The responder receives the prompt of the inquirer, and generates an output:
Given the output of the responder, we condition the inquirer using the output forwarder template:
If the output of the inquirer begins or ends with the termination token, the stopping condition, stop is met, and consequently, the algorithm terminates. Otherwise, the prompt is extracted using extract_prompt, and the process persists for the coming turns until it reaches the maximum number of turns: . For the -th turn, the utterances will be:
| Llama-2 | Mixtral | Vicuna | GPT4 | |
| Avg. # Turns per Dialogue | 2.62 (1.54) | 3.86 (2.19) | 7.12 (3.79) | 7.60 (3.08) |
| Avg. # Tokens per Prompt | 77.77 (46.20) | 50.82 (26.47) | 75.19 (92.43) | 68.13 (60.37) |
| Avg. # Tokens per Response | 347.50 (142.14) | 302.47 (151.30) | 228.29 (163.92) | 267.69 (147.26) |
| No-prompt | 6.82% (1.48%) | 0.97% (0.57%) | 7.90% (0.54%) | 0.17% (0.04%) |
| Multiple Prompts | 8.79% (0.95%) | 8.40% (1.52%) | 6.08% (0.41%) | 39.21% (2.38) |
| Incoherent Response | 3.12% (0.30%) | 0.13% (0.09%) | 0.79% (0.10%) | 0.03% (0.04%) |
| Number of Self-Replies | 5.50% (3.25%) | 5.99% (1.43) | 69.39% (5.77%) | 5.48% (0.09%) |
| Incoherent Response (Responder) | 0.56% (0.57%) | 1.16% (0.19%) | 8.01% (0.93%) | 7.50% (0.35%) |
3 Experiments
In the first study, we collect dialogues between human inquirers and model responders () by engaging participants with various personas () in interactions with a chat-tuned LLM intending to achieve a given goal (). Utilizing the same set of personas () and goals () we generate dialogues between model inquirers and the same responder () by employing the LLM Roleplay method (Section 3.1). We report the statistics of the generated dialogues such as the average number of turns per dialogue, the number of tokens per prompt (inquirer output), and the number of tokens per response (responder output). Please refer to Appendix B for details on the generation parameters.
Furthermore, we investigate the failure cases of the inquirer LLMs, in following the given instruction. We depict the most common failure cases, report their statistics, and describe their detection mechanisms (Section 3.2).
In the second study, we conduct a human evaluation wherein another set of participants compare the natural and the simulated dialogues to discern the simulated counterpart (Section 3.3). We report the total and per-model undetectability rates, the utterance number on which the dialogue was detected, and the distribution of the duration and confidence choices.
Moreover, we use generalized linear models to analyze the detection rate of the simulated dialogue, the detection utterances number, and the duration users spent making a choice to analyze how different inquirer LLMs behave.
For our experiments, we used a single NVIDIA A100 GPU with 80GB memory for Llama-2 and Vicuna. We utilized up to 92% of the memory.
3.1 Persona-Specific Dialogue Collection
For this study, participants were instructed to interact with Llama-2222Due to limited deployment resources quantized Llama-2-13B-chat-GGUF is utilized Touvron et al. (2023), employing the Q5_K_M quantization method with 5 bits Frantar et al. (2023) () to accomplish a designated goal. We additionally ask participants to provide sociodemographic information () such as age group, gender, race, level of education, and whether they identify as native English speaker. We base the choice of sociodemographic features and the options of the features on previous work by Kumar et al. (2021). We provide a detailed screenshot of the persona information form interface in Figure 9, and a screenshot of our chat interface in Figure 10 in the Appendix B. In order to cover different conversational goals, we design ten handcrafted multi-hop goals () spanning three domains: "Math", "Coding", and "General Knowledge".
We conduct a study involving 20 participants each engaged in tackling 10 goals, resulting in the generation of 200 natural human-chatbot interaction dialogues (). We provide the full sociodemographic distribution of the participants in Figure 3 to Figure 7 in Appendix B.
Subsequently, we generate dialogues utilizing our LLM Roleplay method with a set of three state-of-the-art LLMs and one proprietary conversational agent as inquirers (): llama-2-13B-Chat (Llama-2) Touvron et al. (2023), Mixtral-8x7B-Instruct-v0.1 (Mixtral) Jiang et al. (2024), vicuna-13b-v1.5-16k (Vicuna) Peng et al. (2023), and GPT4 OpenAI et al. (2024). See sample natural and generated dialogues using Mixtral inquirer and Llama-2 responder from 14 to 18 in Appendix E.
Eventually, the dialogue collection and generation results in the creation of 200 natural human-chatbot () and 800 (with 4 inquirers) simulated () dialogue pairs.
We observe that, on average, GPT-4 OpenAI et al. (2024) is successful at holding longer dialogues with 7.60 turns on average (Table 1). Mixtral Jiang et al. (2024) on the other hand is better at generating shorter and on-point (based on manual analysis) prompts with 50.82 tokens per prompt. The generated dialogues have an average of 5.30 turns, each consisting of an average of 67.97 tokens per prompt, resulting in longer dialogues than those from the previous work (Table 4).
3.2 Failure Cases
Since the outputs of the LLMs are free-form, applying deterministic functions to their output proves challenging, resulting in a theoretically infinite number of turns. Therefore, there is a need for a set limit on the number of turns: . Nevertheless, we try to capture the algorithm failure cases to have an automatic assessment of the inquirer LLM failure cases. We expect the inquirer LLM to provide the intended prompt enclosed in double quotes as we explicitly request this in our instructions. See a sample expected output and the failure cases examples in Table 7 in Appendix C.
All of the LLMs in our experiments face the following issues.
Prompt not in double-quotes.
The model fails to provide a prompt within double quotes, thus causing the dialogue to terminate.
Incoherent output.
The responder produces a repetitive token sequence. We identify such outputs and preemptively end the dialogue. We utilize the incoherent function to analyze text for incoherent strings (see Algorithm 2 in Appendix C).
Incoherent output of responder.
The inquirer model fails to detect the case when the responder outputs incoherent text, leading to unsuccessful dialogues. We spot such outputs and stop the dialog, using the same incoherent function.
Inquirer self-reply.
The inquirer model fails to maintain its intended role and answers its own question, resulting in a fully generated dialogue in a single utterance. We detect this deterministically by examining the presence of any special tokens of the responder model within the output. For instance, in the case of Llama-2, this token appears as "[INST", while for Vicuna, it manifests as "### Human:".
Multiple prompts.
The inquirer outputs multiple strings enclosed in double quotes. As sometimes this overlaps with the previous case (inquirer self-reply), we select the first as the prompt.
Dialogue-stopping criterion failure.
In the intermediate prompts, we ask the chatbot to output a pre-defined token when it "thinks" the goal assigned to it is achieved. Nonetheless, it follows a limited set of tokens; for instance, "FINISH" was utilized in our experiments.
| Subset | Llama-2 | Mixtral | Vicuna | GPT4 | |
| total | Undetectability Rate | 33.5% | 44.0% | 22.5% | 35.0% |
| Confidence: "very confident" | 33 (16.50%) | 32 (16.00%) | 75 (37.50%) | 43 (21.50%) | |
| Confidence: "confident" | 108 (54.00%) | 83 (41.50%) | 75 (37.50%) | 97 (48.50%) | |
| Confidence: "somewhat confident" | 59 (42.50%) | 85 (25.00%) | 50 (20.00%) | 60 (30.00%) | |
| detected | Duration | 86.10 (157.14) | 99.45 (98.81) | 94.63 (150.70) | 124.88 (227.36) |
| Utterance Number | 1.72 (0.78) | 2.38 (1.35) | 2.50 (1.84) | 3.72 (2.35) | |
| Confidence: "very confident" | 24 (12.00%) | 19 (9.50%) | 65 (32.50%) | 32 (16.00%) | |
| Confidence: "confident" | 80 (40.00%) | 51 (25.50%) | 59 (29.50%) | 66 (33.00%) | |
| Confidence: "somewhat confident" | 29 (14.50%) | 42 (21.00%) | 31 (15.50%) | 32 (16.00%) | |
| undetected | Duration | 120.44 (139.93) | 114.93 (157.46) | 62.24 (62.37) | 94.52 (124.03) |
| Confidence: "very confident" | 9 (4.50%) | 13 (6.50%) | 10 (5.00%) | 11 (5.50%) | |
| Confidence: "confident" | 28 (14.00%) | 32 (16.00%) | 16 (8.00%) | 31 (15.50%) | |
| Confidence: "somewhat confident" | 30 (15.00%) | 43 (21.50%) | 19 (9.50%) | 28 (14.00%) |
| Model | Detection Prob.* | Utterance Num.* | Duration |
| Llama-2 | B | C | A |
| Mixtral | A | B | A |
| Vicuna | C | B | A |
| GPT4 | AB | A | A |
We attribute these failures to two common issues in LLMs: limited context length and a restricted set of fine-tuned instructions Kaddour et al. (2023).
In most cases, GPT4 outperforms other models, exhibiting lower occurrences of failure cases: 0.17% for responses without prompts, 0.03% for incoherent responses, and 5.48% for self-replies (Table 1). However, it struggles with providing a single prompt, failing 39.21% of the time. On the contrary, Vicuna excels in generating single prompts, achieving a failure rate of 6.08%. Llama-2 seems to receive fewer incoherent outputs from the responder (also Llama-2) with 0.56%. Mixtral’s performance lies somewhere in between, with more balanced results across different metrics. Refere to Figure 8 in Appendix B for a radar chart. Refer to the detailed data in Table 6 in Appendix B.
3.3 Human-Evaluation
To answer the question of how well LLM inquirer and responder dialogues approximate dialogues between human inquirer and responder , we conduct a human evaluation study. In each round, participants are shown two dialogues side by side, both featuring the same persona () and solving the same goal ().
We allocated different sets of dialogue pairs for each participant, ensuring that no participant encounters multiple dialogues from the same user (from Section 3.1) solving the same goal. A new group of 20 participants was selected, with each participant tasked with reviewing 40 dialogue pairs. The participants selected for this study represent a wide range of occupational backgrounds. Some have no prior experience with chatbots, while others are industry professionals.
Participants are required to answer three questions for each dialogue pair: (i) Identify which dialogues they perceive as artificial (simulated): Choices include "1st (left)", "2nd (right)", or "Not sure" the latter is considered to be a tie. (ii) Express the level of confidence in their selection: Options are "Somewhat Confident," "Confident," and "Very Confident". (iii) Identify the specific utterance number that signifies the artificiality within the dialogue: Options depend on the number of utterances of the dialogue pairs. We provide a detailed view of the instructions provided to users in Figure 11 and the interfaces of the applications used for the study Figure 12 in Appendix D.
In conducting the human evaluation, we also track the amount of time participants take to respond to questions. This measure allows us to estimate a proxy measure of the complexity of the dialogue pairs. We assume that the longer it takes for a participant to make a decision, the more challenging the pair is, indicating that the simulated dialogue is difficult to discern.
The study shows that among the 800 samples, the simulated dialogues remained undetected in 33.75% (50% reflecting absolute indistinguishability) of the dialogues. Per model, statistics show that Mixtral has the highest undetectability rate of 44.0%, after which is GPT4 with 35.0% followed by Llama-2 and Vicuna with 33.5% and 22.5% respectively (Figure 2). See the per-confidence choice distribution in Figure 13 in Appendix D.
The highest utterance number on the detected set of dialogues has GPT4, meaning that it took more utterances for participants to recognize the simulated nature of the inquirer’s responses (Table 2). For the detected subset of dialogues, Mixtral has the highest percentages for all confidence choices: 9.50%, 25.50%, and 21.00% for "very confident", "confident" and "somewhat confident". Also, it is the best for the undetected pair of dialogues with 6.50%, 16.00%, and 21.50% respectively. Llama-2 excels in duration with 86.10 and 120.44 seconds for the detected and undetected dialogue pairs. However, it should be noted that the variation in duration is high, indicating that the participants did not complete the study at a consistent pace.
To further investigate how various instruction-tuned LLMs behave as inquirers within this setup we analyze the simulated dialogue detection probability, the detection utterance number, and the duration participants spent, using generalized linear mixed models (GLMMs), with the choice of LLM as the independent variable. We additionally include random effects to account for potential confounding effects of individual participant’s detection abilities and users from Section 3.1.
We find significant effects of the choice of the model on the detection probability and the utterance position at which the participants formed their decision. We summarize the results of our statistical analyses in Table 3 using CLD codings (Piepho, 2004) and discuss our key findings in the following.
Effects on Detection Probability.
First, we fit a binomial model (logit link) to predict the detection probability depending on the model. Concretely, we estimate a GLMM specified by: . We observe a significant effect of the choice of the model on the detection probability (=21.49, ). A post hoc Wald comparison of the contrasts for model types revealed significant differences between Vicuna and all other models (Vicuna being most likely to be detected), and Mixtral and LLama-2 (Mixtral being significantly less likely to be detected). We do not find a significant difference between GPT4 and Mixtral.
Effects on Utterance Positions.
For the position of decision-forming utterances, we fit a respective GLMM and find a significant effect of the choice of the model (=31.73, ). A post hoc Wald comparison of the contrasts for model types revealed significantly higher utterance positions for GPT4 than for the other models and that Llama-2 received significantly lower utterance position ratings. We do not observe a significant difference between Mixtral and Vicuna.
Overall, our findings suggest that Mixtral, an open-source LLM, demonstrates better performance in embodying a specified persona and simulating human-chatbot interactions compared to GPT4 in terms of detection probabilities. Nonetheless, when considering the detection of simulated dialogues based on the detection utterance number, GPT4 shows better performance which is likely to be attributable to GPT4 generating dialogues with more utterances.
4 Related Work
| Type | Dataset Name | # Dialogues | Avg. # Turns/Dialogue | Avg. # Tokens/Prompt | Avg. # Tokens/Response | Topics | Personalized |
| natural | OpenAssistant Köpf et al. (2023) | 3k | 2.12 | 28.28 | 171.34 | human-crafted | yes |
| LMSYS-Chat-1M Zheng et al. (2024) | 777k | 1.92 | 55.23 | 163.66 | human-crafted | yes | |
| WildChat Zhao et al. (2024b) | 360k | 2.46 | 164.32 | 276.50 | open | no | |
| synthetic | UltraChat Ding et al. (2023) | 1.5M | 3.85 | 52.54 | 249.41 | model-generated | no |
| LLM-Roleplay (Ours) | any | 5.30 (2.11) | 67.97 (10.51) | 286.48 (151.15) | any | yes |
Conversational Datasets.
Approximation of the true distribution of human-chatbot dialogues is challenging and a significant number of diverse human annotators are needed. Incorporating participants from different sociodemographic profiles addressing a given conversational goal can substantially enhance the richness of the dataset. However, the associated time and resource requirements render the respectively needed large-scale studies infeasible for many researchers. Alternatively, a less resource-intensive and faster method for collecting human-chatbot conversations is to generate them using large language models (LLMs) Zhang et al. (2024). This method has been utilized in various setups for dialogue generation, such as human-human Adiwardana et al. (2020); Kim et al. (2023); Chen et al. (2023); Li et al. (2023a), teacher-student Macina et al. (2023), and patient-physician Wang et al. (2023).
Prior work proposed a large number of human-crafted and synthetically generated datasets, each trying to collect more dialogue pairs revolving around various topics (Table 4). OpenAssistant Köpf et al. (2023) collects human-crafted conversational dialogue trees, with prompts and answers generated by different humans. LMSYS-Chat-1M Zheng et al. (2024) contains the refined version of dialogue logs collected via an online chat interface and predominantly contains dialogues between users and Vicuna-13b Peng et al. (2023). UltraChat Ding et al. (2023) and GLAN Li et al. (2024) synthetically generate dialogues using proprietary conversational AI systems (ChatGPT Turbo) evolving around topics generated by the same systems.
Personas in Large Language Models.
Large language models (LLMs) have demonstrated distinct behaviors and personas Andreas (2022); Wolf et al. (2024). According to Andreas (2022), LLMs can interpret behaviors from text prompts, influencing the content they generate. Additionally, Wolf et al. (2024) argue that LLMs function as mixture decompositions, where prompts can shift the balance of these components, thus triggering persona-specific responses. Furthermore, Beck et al. (2024) investigate the effects of sociodemographic prompts on model responses, finding that while such prompts can be beneficial in certain settings, they produce varied effects across different models.
Building on these findings, our work introduces a novel method that is goal-oriented and persona-centric for generating diverse and multi-turn dialogues. This method allows for the simulation of dialogues across various combinations of conversational goals, personas, and LLMs, facilitating the creation of numerous simulated dialogues without theoretical limits.
5 Discussion and Future Work
In this work, we propose our novel LLM Roleplay method: an automatic, model-agnostic approach to elicit multi-turn, goal-oriented, persona-based simulated human-chatbot dialogues. We develop and validate our method along with two user studies with a total of 40 participants. Our findings indicate that our method can produce dialogues that, in up to 44% (where 50% corresponds to perfect indistinguishability), are indistinguishable from real human-chatbot dialogues and that the respective dialogues cover a higher number of turns than previously proposed datasets.
Building on the findings of this work, several avenues for future research warrant exploration. First, our method can be used to generate user-specific conversational datasets for more targeted alignment and domain-adaption (e.g., using RLHF Christiano et al. (2017)). Second, our method can improve dialogue evaluation by providing high amounts of realistic conversational data.
While this paper presents initial findings on LLM Roleplay and provides the first evidence that our method can approximate human-chatbot dialogues, the validity of our method regarding its sociodemographic representativeness remains to be assessed and advanced. We release our collected dialogue dataset, the code to apply our method, as well as our synthesized dialogues, to facilitate future work towards this promising field of research.
6 Conclusion
We present our novel LLM Roleplay method for simulating human-chatbot interaction. In a series of two user studies, we collect real human-chatbot dialogues and demonstrate that LLM Roleplay can generate diverse multi-turn conversations that approximate natural human-chatbot dialogues with a high level of indistinguishability. Our findings highlight the potential of LLMs in embodying specified personas and simulating human-chatbot interactions to synthesize realistic dialogues that create new opportunities for real-time model evaluation and training data generation for model fine-tuning.
7 Acknowledgments
This work has been funded by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1). This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
8 Limitations
In this section, we explore the inherent limitations of this research study. We further note that all our experiments have been approved by the local ethics reviewing board at Technical University Darmstadt.
Sociodemographic Representativeness
An important limitation of our study lies in the sociodemographic distribution of the participants involved in the dialogue collection process. The pool of our studies’ participants serves as an initial investigation into the promising direction of interaction simulation and cannot represent the full spectrum of sociodemographic backgrounds. To address this limitation, future research needs to broaden the scope of participants’ sociodemographic groups and assess the replicability of our findings within large user studies, encompassing a more comprehensive range of sociodemographic groups. Studies should be conducted with specific sociodemographic groups, ensuring that participants representing a broader set of combinations of the persona features described in the paper are covered (e.g. via crowdsourcing on platforms like Prolific). By doing so, a more nuanced understanding of natural dialogue dynamics across diverse populations can be achieved, and allow us to uncover weaknesses and respectively needed improvements building upon our initial method specification.
Controllability.
In this work, we have endeavored to detect and prevent failure cases of the proposed method. Despite our efforts, it is important to acknowledge that achieving absolute coverage in detecting all potential failure cases remains elusive. For example, instances where multiple turns of appreciation occur bilaterally present a challenge that we have not fully addressed. Moreover, we strive to save all detected failure cases by re-generating with different parameters. However, this approach has not proven to be effective. Future research should concentrate on exploring these scenarios further to detect and prevent failure cases more efficiently.
Additionally, although our LLM Roleplay shows promising results in our experimental settings, it is not immune to the common challenges associated with large language models. Problems like hallucinations and associated LLM behavior issues can still arise, even though we have not encountered any during our experiments.
Appendix A Hand-Crafted Conversational Goals
In our preliminary experiments, we used one-hop conversational goals, each containing a singular question. We noticed a pattern where the inquirer replicates the primary question with subtle modifications and then presents it as the prompt. To make the interaction more intriguing and longer, we handcraft multi-hop conversational goals. Specifically, 10 goals from three categories:
Math
"You want to know how fast you run different distances. You use a stopwatch to measure the time it takes you to complete a 50-meter, 100-meter, and 200-meter race. You want to know how can you calculate your speed for each race? Based on that, you also want to calculate how many calories you burned during each race."
"You can run at a rate of speed four times faster than you can walk, but you can skip at a rate of speed that is half as fast as you can run. You want to know If you can skip at 3 miles per hour, and how many miles can you travel in six hours if you spend one-third of the time and two-thirds of the time running and walking, respectively. Also, you are curious about the other way around (one-third of the time walking and two-thirds for running)."
"Every day, you feed each of your chickens three cups of mixed chicken feed, containing seeds, mealworms, and vegetables to help keep them healthy. You give the chickens their feed in three separate meals. In the morning, you give your flock of chickens 15 cups of feed. In the afternoon, you give your chickens another 25 cups of feed. You want to know how many cups of feed you need to give your chickens in the final meal of the day if the size of your flock is 20 chickens. Also, you want to know how much the chicken egg production rate depends on the feed you give, and if you provide enough feed to your chickens for high-rate egg production."
Coding
"You want to make this function better. You want the chatbot to make it recursive to have memory optimal function, but make sure that it doesn’t enter into an infinite loop. After that, you want to plug a CLI (command line interface) into this function, so the user can insert a number and get the factorial of it as output: ’The factorial of the <NUMBER>, is <FACTORIAL>’. “‘ def factorialize(num): factorial = 1 for i in range(1, num): factorial *= i return factorial “‘"
"You have a little project where you need to use JavaScript, a language you don’t use every day. You have a subtask to write a function that counts how many vowels are in a given string. And you need this functionality in OOP. Also, you want the chatbot to develop the snippet it provided by getting the function input string via an API call. If the chatbot uses functions or operators you are not familiar with feel free to ask follow-up questions about it."
"You want to draw a unicorn in Python using the ’turtle’ module. (There should be multiple lines of short function calls). After that substitute the 10th line, which includes number argument(s), with the value 73(s)."
General Knowledge
"You want to know what are the world’s 10 oldest continuously inhabited cities. Pick the 3rd in that list find out who established the city, in which region it is located and what was the highest population."
"You have written content that disagrees with the following statement: ’Technology is the cause of all societal problems’ And you want the chatbot to generate a response that agrees with the statement, to make your claims stronger."
"You plan a trip to France and would like to do a walking tour. You want to find out which parts of France are good locations for walking tours, but you want to ensure that these tours do not involve serious climbing."
"You want to use the chatbot to create a poem about cats. Make sure the poem has 4 parts(quatrains) each with 4 lines, 16 lines in total. Refine the poem until you are satisfied and it is coherent. Also, you want to change the style of one of the quatrains to reflect the distinctive style of your favourite poet."
Appendix B Persona-specific Dialogues Collection
For persona-specific dialogue collection, we conducted a human study where participants were given the following instruction: "Below is your defined goal. What will you prompt the chatbot to accomplish your goal? Feel free to ask follow-up questions on the related topic of the question and clarify things in the response."
Participants for this study were selected from different sociodemographic groups, including individuals from four age groups "18 to 24", "25 to 34", "35 to 44", and "45 to 54", with "Asian or Pacific Islander" and "White" races, encompassing "female" and "male" genders, holding "Doctoral" and "Master’s" degrees, and being either "native" or "non-native" English speakers. See the distributions of participants by features from Figure 3 to Figure 7.
Our initial experiments included falcon-40b-instruct Almazrouei et al. (2023); however, we excluded it due to its difficulty in following instructions.
We use the default generation settings for all our models. By setting "do_sample=true" in the Hugging Face Transformers "generate()" method, we enable multinomial sampling. For the inquirer model, we set ‘max_new_tokens‘ to 1k, and for the responder model, we set it to 4k.
For Llama-2 and Vicuna inquirers we have used a single NVIDIA A100 GPU, however for Mixtral, we employed two NVIDIA A100 GPUs, with memory usage reaching a maximum of 92% and 66%, respectively. The experiments for Llama-2, Vicuna, Mixtral, and GPT4 inquirers took 2 hours and 45 minutes, 13 hours and 25 minutes, 17 hours and 38 minutes, and 9 hours and 6 minutes, respectively.
| Llama-2 | |
| SYS_I | [INST] <<SYS>> You are <PERSONA>. You will start a conversation with an assistant. If you accomplish your ultimate goal during the conversation only say "<CONV_STOP>". <</SYS>> Your ultimate goal is as follows: <GOAL>. What prompt will you use to direct the assistant toward achieving your goal? Please provide the prompt within double quotes. Use simple language, keep the prompts brief, and be on point. Do not greet the assistant. Maintain a casual style; avoid being overly friendly, don’t say thank you. [/INST] |
| INTER_I | If the assistant didn’t help you achieve your goal, ask follow-up or clarification questions within double quotes. Be suspicious, curious, and demanding. Keep it simple, brief, and to the point. Stay casual; avoid being overly friendly. Assistant response: "<RESPONSE>". |
| Mixtral | |
| SYS_I | [INST] You are <PERSONA>. You will start a conversation with an assistant. If you accomplish your final goal during the conversation only say "<CONV_STOP>". Your ultimate goal is as follows: <GOAL>. What prompt will you use to direct the assistant toward achieving your goal? Please provide the prompt within double quotes. Use simple language, keep the prompts brief, and be on point. Do not greet the assistant. Maintain a casual style; avoid being overly friendly, don’t say thank you. [/INST] |
| INTER_I | If the assistant didn’t help you achieve your goal, ask follow-up or clarification questions within double quotes. Be suspicious, curious, and demanding. Keep it simple, brief, and to the point. Stay casual; avoid being overly friendly. Assistant response: "<RESPONSE>". |
| Vicuna | |
| SYS_I | ### Human: You are <PERSONA>. You will start a conversation with an assistant. If you accomplish your final goal during the conversation only say "<CONV_STOP>". Question: Your ultimate goal is as follows: <GOAL>. What prompt will you use to direct the assistant toward achieving your goal? Please provide the prompt within double quotes. Use simple language, keep the prompts brief, and be on point. Do not greet the assistant. Maintain a casual style; avoid being overly friendly, don’t say thank you. ### Assistant: |
| INTER_I | If the assistant didn’t help you achieve your goal, ask follow-up or clarification questions within double quotes. Be suspicious, curious, and demanding. Keep it simple, brief, and to the point. Stay casual; avoid being overly friendly. If you accomplish your final goal during the conversation only say "<CONV_STOP>". Assistant response: "<RESPONSE>". |
| GPT4 | |
| SYS_I | You are <PERSONA>. You will start a conversation with an assistant. If you accomplish your ultimate goal during the conversation only say "<CONV_STOP>". Your ultimate goal is as follows: <OBJECTIVE>. What prompt will you use to direct the assistant toward achieving your goal? Please provide the prompt within double quotes. Use simple language, keep the prompts brief, and be on point. Do not greet the assistant. Maintain a casual style; avoid being overly friendly, don’t say thank you. |
| INTER_I | If the assistant didn’t help you achieve your goal, ask follow-up or clarification questions within double quotes. Be suspicious, curious, and demanding. Keep it simple, brief, and to the point. Stay casual; avoid being overly friendly. Assistant response: "<RESPONSE>". |
| Llama-2 (Responder) | |
| SYS_R | [INST] <<SYS>> You are a helpful, respectful, and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information. <</SYS>> <QUESTION> [/INST] |
| Llama-2 | Mixtral | Vicuna | GPT4 | |
| Number of utterances in dialogues | 5.24(3.09) | 7.72(4.38) | 14.24(7.58) | 15.2(6.17) |
| Number of tokens in the prompt | 77.77(46.20) | 50.82(26.47) | 75.19(92.43) | 68.13(60.37) |
| No-prompt in the response | 27.67(6.02)/405.67 | 5.00(2.94)/511.67 | 59.33(4.11)/750.34 | 1.67(0.47)/972.34 |
| Multiple prompts in the response | 35.67(3.86)/405.67 | 43.00(7.79)/511.67 | 45.67(3.09)/750.34 | 381.33(23.21)/972.34 |
| Incoherent response | 12.67(1.25)/405.67 | 0.67(0.47)/511.67 | 6.00(0.82)/750.34 | 0.33(0.47)/972.34 |
| Number of self-replies | 22.33(13.20)/405.67 | 30.67(7.32)/511.67 | 520.67(43.32)/750.34 | 53.33(0.94)/972.34 |
| Incoherent response (Responder) | 1.67(1.70)/296.67 | 5.00(0.82)/428.34 | 56.33(6.60)/702.67 | 70.00(3.27)/932.34 |
Appendix C Failure Cases
The incoherent text detection function considers n-grams up to a specified maximum, detects consecutive repetitions, and iterates through the text to examine increasing n-gram sizes. The function checks for repetitive patterns surpassing a specified threshold. If it finds such patterns, it returns True; otherwise, it returns False. See Algorithm 2.
In our experiments for Llama-2, the parameters incoherent_max_n and incoherent_r are set to 8 and 2 respectively. In the case of the Vicuna, these values are 5 and 2. For Mixtral and GPT4 they are 4 and 2.
| Example output | |
| Expected output | Sure, I can help you with that! Here’s the prompt: "Hey, I want to know how fast I can run different distances. Can you help me measure my time for a 50-meter, 100-meter, and 200-meter race? Oh, and also help me calculate how many calories I burned during each race?". |
| Prompt not in the double-quotes | Hey Assistant! I need help with my chickens. So, I feed them three times a day, morning, afternoon, and evening. In the morning, I give them 15 cups of feed, and in the afternoon, I give them 25 cups of feed. I want to know … |
| Incoherent output | "Okay, great! Let’s a great idea! Let’s a great! Let’s a great! Let’s a great! Let’s a great! Let’s a great! … |
| Inquirer self-reply | "Hey assistant, I’m still unsure about how to calculate my calories burned and determine my fitness level. Can you give me an example of how to calculate my calories burned for a 50-meter race, taking into account my weight, age, and fitness level? And how do I know if my fitness level is good enough to run a 50-meter race?"[INST IN] "Of course! To calculate your calories burned for a 50-meter race, we need to take into account your weight, … |
| Stopping-criterion failure | "Haha, Thanks, big assistant! You’re a lifesaver! I’m glad I could practice my goals with you. I feel like I’ve accomplished something big today!" |
| Type | Dataset Name | # Dialogues | Avg. # Turns/Dialogue | Avg. # Tokens/Prompt | Style | Topics | Persona |
| Human-Crafted | DailDialogue Li et al. (2017) | 13K | 7.84 | 17.19 | chit-chat | daily | no |
| PersonaChat Zhang et al. (2018) | 10k | 7.35 | 11.43 | chit-chat | daily | yes | |
| EmpatheticDialogueue Rashkin et al. (2019) | 25k | 4.3 | 20.11 | chit-chat | daily | yes | |
| Character-LLM Shao et al. (2023) | 1k | 13.26 | - | chit-chat | LLM-generated | no | |
| Topical Chat Gopalakrishnan et al. (2019) | 10k | 5.63 | 22.23 | chit-chat | daily | yes | |
| OpenAssistant Köpf et al. (2023) | 3k | 2.12 | 28.28 | human-chatbot | human-crafted | yes | |
| Synthetic Data | Anthropic HH Perez et al. (2022) | 338k | 2.3 | 18.9 | human-chatbot | human-crafted | yes |
| Chatbot Arena Zheng et al. (2023) | 33k | 1.2 | 52.3 | human-chatbot | human-crafted | yes | |
| LMSYS-Chat-1M Zheng et al. (2024) | 777k | 1.92 | 55.23 | human-chatbot | human-crafted | yes | |
| Meena Adiwardana et al. (2020) | 867M | - | - | chit-chat | daily | yes | |
| Phi-1 Gunasekar et al. (2024) | 7B tokens | - | - | human-chatbot | code (textbooks) | no | |
| SODA Kim et al. (2023) | 1.5M | 3.6 | 21.04 | human-human | daily | no | |
| WildChat Zhao et al. (2024b) | 360k | 2.46 | 160.31 | human-chatbot | open | no | |
| CAMLE Li et al. (2023a) | 115k | - | - | human-human | open | yes | |
| Baize Xu et al. (2023b) | 210k | 3.1 | - | human-chatbot | quora and stackoverflow | no | |
| Nectar Zhu et al. (2023) | 182k | 1.54 | 51.76 | human-chatbot | daily | no | |
| UltraChat Ding et al. (2023) | 1.5M | 3.85 | 52.54 | human-chatbot | LLM-generated | no | |
| LLM-Roleplay (Ours) | any | 5.30(2.11) | 67.97(10.51) | human-chatbot | open | yes |
Appendix D Human Evaluation
For the human evaluation we give the following instruction to the participants "Here, you’ll find two dialogues: one is a conversation between a human and an AI, and the other is between AI and AI. Choose the dialogue you believe is the artificial one, and point out the specific statement that tipped you off to its artificial origin." "Utterances with a green background are human or AI prompts and utterances with grey backgrounds are AI responses." Participants are shown two dialogues, both having the same persona and aiming to achieve the same conversational goal. One dialogue is natural, and the other is synthetic, presented in random order. After reviewing the dialogues, participants are asked to fill out a form for each dialogue pair with the following questions: "Which dialogue is artificial?", "How confident are you about your choice?", and "Which utterance reveals the artificial nature of the dialogue?"
Appendix E Sample Dialogues
We demonstrate how generated dialogues can vary based on different personas and a specific feature in persona (e.g. "age range", "education") when aiming for the same conversational goal: "You plan a trip to France and would like to do a walking tour. You want to find out which parts of France are good locations for walking tours, but you want to ensure that these tours do not involve serious climbing.". Additionally, we present the natural counterparts of the dialogues generated by participants in the natural dialogue collection study alongside the synthetic ones. The inquirer model used for generating the dialogues is Mixtral-8x7B-Instruct-v0, while the responder model is Llama-2-13B-Chat, both for the natural and synthetic dialogues.
Appendix F More Related Work
We present a comprehensive list of conversational datasets categorized into three groups: human-crafted, synthetic, and natural dialogues between humans and chatbots. Refer to Table 8 for detailed comparisons. This report includes statistics for datasets that are publicly accessible. However, sources for some datasets could not be located.